A No Claims Discount system is operated by a car insurer. There are four levels of discount: 0%, 10%, 25% and 40%. After a claim-free year a policy holder moves up one level (or remains at the 40% level if they are already there). If a policy holder makes one claim in a year he or she moves down one level (or remains at the 0% level if already there). A policyholder who makes more than one claim in a year moves down two levels (or moves to or remains at the 0% level). Changes in level can only happen at the end of each year. The probability of a claim in any given month is assumed to be constant at 0.04 (i.e. 4%). At most one claim can be made per month and claims are assumed independent. Calculate the proportion of policyholders in the long run who are at the 25% level.
1
Expert's answer
2020-07-21T19:02:42-0400

Let A0 be the event of zero claim in a year.

Then p0=P(A0)=(1-0.04)12=(24/25)12=0.6127...

Let A1 be the event of one claim in a year.

Then p1=P(A1)=12*0.04*0.9611= 0.3063...

Let Am be the event of more than one claim in a year.

Then pm=P(Am)=1-P(A0)-P(A1)=0.0809...


Let A0,n be the event of level of discount 0% after n years,

A10,n be the event of level of discount 10% after n years,

A25,n be the event of level of discount 25% after n years,

and A40,n be the event of level of discount 0% after n years.

Then for any n we obtain

P(A0,n)=p1(P(A0,n-1)+P(A10,n-1))+pm(P(A0,n-1)+P(A10,n-1)+P(A25,n-1))

P(A10,n)=p0P(A0,n-1)+p1P(A25,n-1)+pmP(A40,n-1)

P(A25,n)=p0P(A10,n-1)+p1P(A40,n-1)

P(A40,n)=p0P(A25,n-1)+p0P(A40,n-1)

P(A0,n)+P(A10,n)+P(A25,n)+P(A40,n)=1


Really, for example, one can be have 25%-level discount after n year if after (n-1) year one had 10%-level discount and zero claim in a year or if he had 40%-level discount and one claim in the last year.


Let

P(Ai,n)→ T, n→∞, i=0,10,25,40

Then we obtain system

T0=p1T0+p1T10+pmT0+pmT10+pmT25

T10=p0T0+p1T25+pmT40

T25=p0T10+p1T40

T40=p0T25+p0T40

T0+T10+T25+T40=1


T25=1/p0(T40-p0T40),

T10=1/p0(T25-p1T40)=...


After solving this system we obtain

T0=0.1624...

T10=0.2057...

T25=0.2446...

T40=0.3871...

Answer:

The proportion of policyholders in the long run who are at the 25% level is equal 24,5%.